# A Fourier-based inference method for learning interaction kernels in particle systems

### v1

@@ 1 Introduction
Stochastic interacting particle systems find applications in many areas related to social sciences [18], collective behavior [35], pedestrian dynamics [24], physics [22], biology [46], and machine learning [45]. Recently, tremendous progress has been made in the qualitative and quantitative understanding of such systems, from different perspectives: modeling, analysis, quantitative propagation of chaos [28, 15], optimal control [8], as well as the development of numerical methods for solving the nonlinear, nonlocal SDEs and PDEs that we obtain in the mean-field limit [23, 11, 30]. We refer to [9, 10] for a recent review and references to the literature. In addition, several important contributions have been made to the development of efficient inference methodologies for interacting particle systems and their mean-field limit. Statistical inference, learning, data assimilation, and inverse problems for McKean SDEs and McKean–Vlasov PDEs are topics of great current interest. In particular, the wealth of data in real-world problems pushes towards data-driven models, and therefore learning parameters or functions in mathematical models from observed data is fundamental. A partial list of recent activities on inference for mean-field SDEs and PDEs includes kernel methods [31, 29], constrast functions based on a pseudo-likelihood [3], maximum likelihood estimation [17], stochastic gradient descent [38, 44], approximate likelihood based on an empirical approximation of the invariant measure [19], method of moments [40], and eigenfunction martingale estimating functions [39, 41]. An important observation upon which our previous work on inference for McKean SDEs is based is that, under the assumption of uniform propagation of chaos, the nonlinear mean-field SDE can be replaced by a linear (in the sense of Mckean) SDE, obtained by calculating the convolution term in the drift with respect to the (unique) invariant measure of the process. A detailed analysis of this approach can be found in [41]. Furthermore, a fully nonparametric estimation of the interaction potential has been studied using different methodologies based on deconvolution [2], data-driven kernel estimators [16], the method of moments [13], least squares [29], and projection techniques [12]. Nonparametric inference for the McKean–Vlasov PDE can also be formulated as an inverse PDE problem and a Bayesian approach can be applied [37].
@@
where $t\in[0,T]$, $V,W\colon\mathbb{R}\to\mathbb{R}$ are the confining and interaction potentials, respectively, and $\sigma>0$ is the diffusion coefficient. Moreover, $\{B_{t}^{(n)}\}_{n=1}^{N}$ are standard independent one-dimensional Brownian motions. We note that the method developed in this paper can be, in principle, extended to the multidimensional case, as we discuss in Section 5. We assume chaotic initial conditions with distribution $\nu$, which is, of course, independent of the Brownian motions $\{B_{t}^{(n)}\}_{n=1}^{N}$. Then, in the limit as $N\to\infty$, and under appropriate assumptions on the confining and interaction potentials, each particle converges in law to the solution of the mean-field McKean SDE
@@
As is well known, without convexity assumptions on interaction and / or confining potentials, the McKean SDE (1.2) can have multiple stationary states [14, 48]. However, in this work, we place ourselves in the (strict) convexity/uniform propagation of chaos setting [32, 28] that ensures the uniqueness of stationary states for the McKean-Vlasov dynamics. We also need the initial distribution to have bounded moments of any order. In particular, we make the following assumption.
@@
Assumption 1.1 is satisfied, for example, in the setting of [32], where the confining and interaction potentials are strictly convex and convex, respectively. However, we believe that uniform propagation of chaos is not needed for the inference methodology presented in this work and that the analysis can be extended to the setting considered in [34]. This is confirmed in the numerical experiments that are presented in Section 4. The analysis of our inference methodology in the presence of phase transitions will be presented elsewhere.
In this work, we consider the problem of learning the interaction kernel $W^{\prime}$ (or equivalently the interaction potential $W$) by observing a single particle from the interacting particle system (1.1). The confining potential is assumed to be known. It is important to note that it is not possible, in general, to identify both the confining and interaction potentials from a single-particle observation. This was already noted in our previous work [39, 40]. In this paper, we make this remark more precise; see Proposition 3.4. We can consider the case where we either observe a continuous trajectory $(Y_{t})_{t\in[0,T]}$ or discrete-time samples from it $\{\widetilde{Y}_{i}\}_{i=0}^{I}$, where
@@
for a particle $\bar{n}$ in (1.1) and sampling rate $\Delta=T/I$. Clearly, because of the exchangeability of the interacting particle system, it does not matter which particle we are observing. Assuming now that $W^{\prime}\in L^{2}(\rho)$ with $\rho$ the unique invariant measure of the mean-field system, we consider the truncated generalized Fourier series expansion
@@
where $\{\psi_{k}\}_{k=0}^{\infty}$ are orthogonal polynomials with respect to the invariant measure $\rho$ that, in practice, will be approximated using the available observations. We then propose to estimate the coefficients $\{\beta_{k}\}_{k=0}^{K}$ in the expansion by solving a linear system that is obtained by imposing appropriate constraints on the moments of $\rho$. These conditions, in turn, are derived from the stationary Fokker–Planck equation (1.4). An outline of the main steps of this approach is provided in Algorithm 1 in Section 3. We note here the link between this part of the proposed statistical inference methodology and the problem of identifying all generators of reversible diffusions–or, equivalently, all Gibbs measures–whose eigenfunctions are orthogonal polynomials with respect to the invariant measure; see [5, Section 2.7] and, in particular, [6] for details. Our method builds upon and connects our two previous papers [39, 13] on the application of the method of moments and of eigenfunction expansions to inference problems for interacting particle systems, respectively. In [39] the main limitation is that the drift, interaction, and diffusion functions are assumed to be polynomials, and, therefore, the inference problem is fully parametric. In an interesting recent paper [13], this hypothesis is eliminated, and no assumptions are made about the functional form of potentials. However, for their methodology to work, multiple stationary particle trajectories, for example, 4, of the mean-field SDE must be observed. In addition, the basis functions in the generalized Fourier series expansion in (1.8) have to be chosen appropriately a priori. The two main innovations of the work reported in this paper are: (1) we only need to observe a single non-stationary particle trajectory of the interacting particle system and not of the mean-field SDE; and (2) we consider basis functions that are purpose-built and tailored to the stochastic interacting particle system. These basis functions are orthogonal polynomials with respect to the invariant measure of the mean-field dynamics.
@@
The rest of the paper is organized as follows. In Section 2 we consider orthogonal polynomials with respect to the invariant measure of the mean-field dynamics, which we approximate using the available observations from a single interacting particle, and in Section 3 we present a generalized method of moments to infer the interaction kernel. The convergence analysis in the limit of infinite data and particles for both orthogonal polynomials and the kernel estimator is provided in Sections 2.1 and 3.2, respectively. Finally, we present different numerical experiments to test our method in Section 4, and we suggest possible developments in Section 5.
@@
$\psi_{k}(x)=\frac{1}{c_{k}}\det\left(\begin{bmatrix}\mathbb{M}^{(0)}&\mathbb{M}^{(1)}&\mathbb{M}^{(2)}&\cdots&\mathbb{M}^{(k)}\\ \mathbb{M}^{(1)}&\mathbb{M}^{(2)}&\mathbb{M}^{(3)}&\cdots&\mathbb{M}^{(k+1)}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ \mathbb{M}^{(k-1)}&\mathbb{M}^{(k)}&\mathbb{M}^{(k+1)}&\cdots&\mathbb{M}^{(2k-1)}\\ 1&x&x^{2}&\cdots&x^{k}\end{bmatrix}\right)\eqqcolon\frac{1}{c_{k}}\det(\mathcal{M}_{k}),$ (2.3)
@@
$\lambda_{kj}$ $=(-1)^{k+j}\det\left(\begin{bmatrix}\mathbb{M}^{(0)}&\cdots&\mathbb{M}^{(j-1)}&\mathbb{M}^{(j+1)}&\cdots&\mathbb{M}^{(k)}\\ \mathbb{M}^{(1)}&\cdots&\mathbb{M}^{(j)}&\mathbb{M}^{(j+2)}&\cdots&\mathbb{M}^{(k+1)}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ \mathbb{M}^{(k-1)}&\cdots&\mathbb{M}^{(k+j-2)}&\mathbb{M}^{(k+j)}&\cdots&\mathbb{M}^{(2k-1)}\end{bmatrix}\right)$ (2.5)
@@
$\widetilde{\mathbb{M}}^{(r)}_{T,N}=\frac{1}{T}\int_{0}^{T}Y_{t}^{r}\,\mathrm{d}t,\qquad\widetilde{\mathbb{M}}^{(r)}_{T,I}=\frac{1}{I}\sum_{i=1}^{I}\widetilde{Y}_{i}^{r},$ (2.7)
depending on whether we have continuous-time or discrete-time observations. We can then build an approximation $\{\widetilde{\psi}_{k}\}_{k=0}^{\infty}$ of the orthonormal basis using formula (2.3), where the exact moments are replaced by the empirical moments. Notice that we denote by $\widetilde{\lambda}_{kj}$ and $\widetilde{c}_{k}$ the coefficients and the normalization constant of the polynomial $\widetilde{\psi}_{k}$, respectively.
@@
We finally recall the result from [40, Lemma 4.3], which, due to Assumption 1.1, using ergodicity and propagation of chaos, and assuming the particles to be initially distributed accordingly to the invariant measure of the mean-field dynamics, i.e. $\nu(\mathrm{d}x)=\rho(x)\,\mathrm{d}x$, quantifies the error given by the approximation of the moments. In particular, there exists a constant $\widetilde{C}>0$ independent of $T$ and $N$, such that for all $q\in[1,2)$
@@
and this estimate will be useful throughout the paper. We note that, as explained in [40, Remark 4.4], the stationarity assumption is made only to simplify the analysis. Due to geometric ergodicity and uniform propagation of chaos, it should be sufficient to start either at a deterministic initial condition or at a distribution that has, for example, finite relative entropy with respect to the invariant measure and finite moments of all orders. In fact, in all the numerical experiments in Section 4, we do not start at stationarity, and the particle trajectory that we observe and that we use to estimate the moments is not stationary.
@@ 2.1 Convergence analysis for the approximated orthogonal polynomials
Under Assumption 1.1, for all $k,j\geq 0$ and for all $q\in[1,2)$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
First, by [27, Theorem 2.12], we have
@@
where we remark that $q/2+1\in[3/2,2)$. Then, since the spectral norm is bounded by the Frobenius norm and due to the uniform boundedness of the moments and the estimate (2.8), following the proof of [40, Lemma 4.5(i)] we obtain the desired result. We finally remark that the dependence of $C$ on $j$ is not explicitly emphasized since $j\leq k$. ∎
@@
Under Assumption 1.1, for all $k\geq 0$ and for all $q\in[1,2)$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
Then, applying Hölder’s inequality, the boundedness of the moments, the estimates (2.8), and Lemma 2.2, we obtain (i). Notice now that we have
@@
Notice that Lemma 2.3 gives bounds on the difference between the normalization constants. However, $c_{k}$ and $\widetilde{c}_{k}$ appear in the denominator of orthogonal polynomials. Therefore, we need to obtain an estimate in $L^{2}$ of the difference of their inverse, for which the following condition on the boundedness of negative moments for $\widetilde{c}_{k}$ is required.
@@
Assumption 2.4 is always satisfied as long as $\widetilde{c}_{k}$ is bounded away from zero independently of $T$ and $N$, e.g. $\widetilde{c}_{k}\geq\zeta c_{k}$ for some $\zeta>0$. Moreover, in all the numerical experiments that we performed, we observed that Assumption 2.4 holds in practice. We remark that, without this condition, the convergence analysis presented in the following still goes through. However, we could only achieve convergence in probability and without quantitative convergence rates.
@@
Under Assumptions 1.1 and 2.4, for all $k\geq 0$ and for all $q\in[1,2r/(r+2))$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
Using Hölder’s inequality with exponents $r/q$ and $r/(r-q)$, we obtain
@@
The desired result then follows from Assumption 2.4 and Lemma 2.3(ii). ∎
@@
Under Assumptions 1.1 and 2.4, for all $k\geq 0$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
which, due to Assumptions 2.4, 2.2, and 2.6, yields the desired result. ∎

### v2

@@ 1 Introduction
Stochastic interacting particle systems find applications in many areas related to social sciences [19], collective behavior [39], pedestrian dynamics [25], physics [23], biology [51], and machine learning [50]. Recently, tremendous progress has been made in the qualitative and quantitative understanding of such systems, from different perspectives: modeling, analysis, quantitative propagation of chaos [29, 16], optimal control [8], as well as the development of numerical methods for solving the nonlinear, nonlocal SDEs and PDEs that we obtain in the mean-field limit [24, 12, 32]. We refer to [10, 11] for a recent review and references to the literature. In addition, several important contributions have been made to the development of efficient inference methodologies for interacting particle systems and their mean-field limit. Statistical inference, learning, data assimilation, and inverse problems for McKean SDEs and McKean–Vlasov PDEs are topics of great current interest. In particular, the wealth of data in real-world problems pushes towards data-driven models, and therefore learning parameters or functions in mathematical models from observed data is fundamental. A partial list of recent activities on inference for mean-field SDEs and PDEs includes kernel methods [33, 30], constrast functions based on a pseudo-likelihood [3], maximum likelihood estimation [34, 18], stochastic gradient descent [43, 49], approximate likelihood based on an empirical approximation of the invariant measure [20], method of moments [45], eigenfunction martingale estimating functions [44, 46], and regularized variational approaches [9]. An important observation upon which our previous work on inference for McKean SDEs is based is that, under the assumption of uniform propagation of chaos, the nonlinear mean-field SDE can be replaced by a linear (in the sense of Mckean) SDE, obtained by calculating the convolution term in the drift with respect to the (unique) invariant measure of the process. A detailed analysis of this approach can be found in [46]. Furthermore, a fully nonparametric estimation of the interaction potential has been studied using different methodologies based on deconvolution [2], data-driven kernel estimators [17], the method of moments [14], least squares [30], and projection techniques [13]. Nonparametric inference for the McKean–Vlasov PDE can also be formulated as an inverse PDE problem and a Bayesian approach can be applied [41]. We also mention the work [35], where a nonparametric least squares estimator that does not suffer from the curse of dimensionality is proposed to learn interaction kernels in deterministic dynamical systems.
@@
where $t\in[0,T]$, $V,W\colon\mathbb{R}\to\mathbb{R}$ are the confining and interaction potentials, respectively, and $\sigma>0$ is the diffusion coefficient. Moreover, $\{B_{t}^{(n)}\}_{n=1}^{N}$ are standard independent one-dimensional Brownian motions. We note that the method developed in this paper can be, in principle, extended to the multidimensional case, as we discuss in Section˜5. We assume chaotic initial conditions with distribution $\nu$, which is, of course, independent of the Brownian motions $\{B_{t}^{(n)}\}_{n=1}^{N}$. Then, in the limit as $N\to\infty$, and under appropriate assumptions on the confining and interaction potentials, each particle converges in law to the solution of the mean-field McKean SDE
@@
A diagram illustrating the relationships between the equations introduced so far is shown in Figure˜1. As is well known, without convexity assumptions on interaction and/or confining potentials, the McKean SDE (1.2) can have multiple stationary states [15, 53]. However, in this work, we place ourselves in the (strict) convexity/uniform propagation of chaos setting [36, 29] that ensures the uniqueness of stationary states for the McKean-Vlasov dynamics. We also need the initial distribution to have bounded moments of any order. In particular, we make the following assumption.
@@
Assumption˜1.1 is satisfied, for example, in the setting of [36], where the confining and interaction potentials are strictly convex and convex, respectively. However, we believe that uniform propagation of chaos is not needed for the inference methodology presented in this work and that the analysis can be extended to the setting considered in [38]. This is confirmed in the numerical experiments that are presented in Section˜4. The analysis of our inference methodology in the presence of phase transitions will be presented elsewhere.
[…]
Figure 1: Schematic illustration summarizing the key mathematical objects and their relationships.
In this work, we consider the problem of learning the interaction kernel $W^{\prime}$ (or equivalently the interaction potential $W$) by observing a single particle from the interacting particle system (1.1). The confining potential is assumed to be known. Intuitively, the approach that we present requires only a single trajectory because it is based on the stationary Fokker–Planck equation (1.4). Due to uniform propagation of chaos, all particles become indistinguishable as $N$ grows large. Moreover, the system converges exponentially fast to equilibrium, so for large $T$, the law of the particle is close to the unique invariant measure. Identifiability conditions for the interaction kernel, when there is no confining potential and the observation is of the mean-field PDE solution, have been studied in [31]. However, it is important to note that it is not possible, in general, to identify both the confining and interaction potentials from a single-particle observation. This was already noted in our previous work [44, 45]. In this paper, we make this remark more precise; see Proposition˜3.5. We can consider the case where we either observe a continuous trajectory $(Y_{t})_{t\in[0,T]}$ or discrete-time samples from it $\{\widetilde{Y}_{i}\}_{i=0}^{I}$, where
@@
for a particle $\bar{n}$ in (1.1) and sampling rate $\Delta=T/I$ fixed (low-frequency regime). Clearly, because of the exchangeability of the interacting particle system, it does not matter which particle we are observing. Assuming now that $W^{\prime}\in L^{2}(\rho)$ with $\rho$ the unique invariant measure of the mean-field system, we consider the truncated generalized Fourier series expansion
@@
where $\{\psi_{k}\}_{k=0}^{\infty}$ are orthogonal polynomials with respect to the invariant measure $\rho$ that, in practice, will be approximated using the available observations. We then propose to estimate the coefficients $\{\beta_{k}\}_{k=0}^{K}$ in the expansion by solving a linear system that is obtained by imposing appropriate constraints on the moments of $\rho$. These conditions, in turn, are derived from the stationary Fokker–Planck equation (1.4). An outline of the main steps of this approach is provided in Algorithm˜1 in Section˜3. We note here the link between this part of the proposed statistical inference methodology and the problem of identifying all generators of reversible diffusions–or, equivalently, all Gibbs measures–whose eigenfunctions are orthogonal polynomials with respect to the invariant measure; see [5, Section 2.7] and, in particular, [6] for details. Our method builds upon and connects our two previous papers [44, 14] on the application of the method of moments and of eigenfunction expansions to inference problems for interacting particle systems, respectively. In [44] the main limitation is that the drift, interaction, and diffusion functions are assumed to be polynomials, and, therefore, the inference problem is fully parametric. In an interesting recent paper [14], this hypothesis is eliminated, and no assumptions are made about the functional form of potentials. However, for their methodology to work, multiple stationary particle trajectories, for example, 4, of the mean-field SDE must be observed. In addition, the basis functions in the generalized Fourier series expansion in (1.8) have to be chosen appropriately a priori. The two main innovations of the work reported in this paper are: (1) we only need to observe a single non-stationary particle trajectory of the interacting particle system and not of the mean-field SDE; and (2) we consider basis functions that are purpose-built and tailored to the stochastic interacting particle system. These basis functions are orthogonal polynomials with respect to the invariant measure of the mean-field dynamics.
@@
The rest of the paper is organized as follows. In Section˜2 we consider orthogonal polynomials with respect to the invariant measure of the mean-field dynamics, which we approximate using the available observations from a single interacting particle, and in Section˜3 we present a generalized method of moments to infer the interaction kernel. The convergence analysis in the limit of infinite data and particles for both orthogonal polynomials and the kernel estimator is provided in Sections˜2.1 and 3.2, respectively. Finally, we present different numerical experiments to test our method in Section˜4, and we suggest possible developments in Section˜5.
@@
[…]
@@
[…]
@@
$\widetilde{\mathbb{M}}^{(r)}_{T,N}=\frac{1}{T}\int_{0}^{T}Y_{t}^{r}\,\mathrm{d}t,\qquad\widetilde{\mathbb{M}}^{(r)}_{I,N}=\frac{1}{I}\sum_{i=1}^{I}\widetilde{Y}_{i}^{r},$ (2.7)
depending on whether we have continuous-time or discrete-time observations. We remark that the subscript $N$ indicates that the empirical moments depend (implicitly) on the number of particles in the system, as the observations are obtained from an interacting particle system of size $N$. We can then build an approximation $\{\widetilde{\psi}_{k}\}_{k=0}^{\infty}$ of the orthonormal basis using formula (2.3), where the exact moments are replaced by the empirical moments. Notice that we denote by $\widetilde{\lambda}_{kj}$ and $\widetilde{c}_{k}$ the coefficients and the normalization constant of the polynomial $\widetilde{\psi}_{k}$, respectively.
@@
We finally recall the result from [45, Lemma 4.3], which, due to Assumption˜1.1, using ergodicity and propagation of chaos, and assuming the particles to be initially distributed accordingly to the invariant measure of the mean-field dynamics, i.e. $\nu(\mathrm{d}x)=\rho(x)\,\mathrm{d}x$, quantifies the error given by the approximation of the moments. In particular, there exists a constant $\widetilde{C}>0$ independent of $T$ and $N$, such that for all $q\in[1,2)$
@@
and this estimate will be useful throughout the paper. We note that, as explained in [45, Remark 4.4], the stationarity assumption is made only to simplify the analysis. Due to geometric ergodicity and uniform propagation of chaos, it should be sufficient to start either at a deterministic initial condition or at a distribution that has, for example, finite relative entropy with respect to the invariant measure and finite moments of all orders. In fact, in all the numerical experiments in Section˜4, we do not start at stationarity, and the particle trajectory that we observe and that we use to estimate the moments is not stationary.
@@ 2.1 Convergence analysis for the approximated orthogonal polynomials
Let $\lambda_{kj}$ be defined in equation (2.5), and let $\widetilde{\lambda}_{kj}$ be its approximation. Under Assumption˜1.1, for all $k,j\geq 0$ and for all $q\in[1,2)$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
First, by [28, Theorem 2.12], we have
@@
where we remark that $q/2+1\in[3/2,2)$. Then, since the spectral norm is bounded by the Frobenius norm and due to the uniform boundedness of the moments and the estimate (2.8), following the proof of [45, Lemma 4.5(i)] we obtain the desired result. We finally remark that the dependence of $C$ on $j$ is not explicitly emphasized since $j\leq k$. ∎
@@
Let $c_{k}$ be defined in equation (2.6), and let $\widetilde{c}_{k}$ be its approximation. Under Assumption˜1.1, for all $k\geq 0$ and for all $q\in[1,2)$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
Then, applying Hölder’s inequality, the boundedness of the moments, estimate (2.8), and Lemma˜2.2, we obtain (i). Notice now that we have
@@
Notice that Lemma˜2.3 gives bounds on the difference between the normalization constants. However, $c_{k}$ and $\widetilde{c}_{k}$ appear in the denominator of orthogonal polynomials. Therefore, we need to obtain an estimate in $L^{2}$ of the difference of their inverse, for which the following condition on the boundedness of negative moments for $\widetilde{c}_{k}$ is required.
@@
Assumption˜2.4 is always satisfied as long as $\widetilde{c}_{k}$ is bounded away from zero independently of $T$ and $N$, e.g. $\widetilde{c}_{k}\geq\zeta c_{k}$ for some $\zeta>0$. Moreover, in all the numerical experiments that we performed, we observed that Assumption˜2.4 holds in practice. We remark that, without this condition, the convergence analysis presented in the following still goes through. However, we could only achieve convergence in probability and without quantitative convergence rates.
@@
Let $c_{k}$ be defined in equation (2.6), and let $\widetilde{c}_{k}$ be its approximation. Under Assumptions˜1.1 and 2.4, for all $k\geq 0$ and for all $q\in[1,2r/(r+2))$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
Using Hölder’s inequality with exponents $r/q$ and r/(rq)r/(r-q), we obtain
@@
The desired result then follows from Assumption˜2.4 and Lemma˜2.3(ii). ∎
@@
Let $\psi_{k}$ be defined in equation (2.4), and let $\widetilde{\psi}_{k}$ be its approximation. Under Assumptions˜1.1 and 2.4, for all $k\geq 0$ there exists a constant $C=C(k)>0$, independent of $T$ and $N$, such that
@@
which, due to Assumptions˜2.4, 2.2, and 2.6, yields the desired result. ∎